# Think How to Think: Mitigating Overthinking with Autonomous Difficulty Cognition in Large Reasoning Models

### v2

@@ Conclusion
This paper presents TH2T, a novel two-stage fine-tuning strategy for large reasoning models (LRMs) that effectively addresses the challenge of overthinking. Our empirical analysis reveals that current LRMs are primarily limited to recognizing task difficulty, exhibiting a necessity for aligning with human System 1 and System 2 thinking patterns. Motivated by this, TH2T progressively enables native difficulty and redundancy cognition based on a heterogeneous dataset and proposed self-hypnosis mechanism. To the best of our knowledge, we are the first to inject difficulty/redundancy/looping hypnosises as built-in triggers for reasoning trajectory intervention. Experiments demonstrate that TH2T significantly reduces inference length with maintained performance stability, exhibiting clear difficulty and redundancy sensitivity. Future research can explore finer-grained reasoning trajectory intervention.
Supplementary Materials
More Experimental Details
Dataset Construction
The datasets used in this paper incorporate difficulty level labeling as a key consideration, so we mainly use mathematical reasoning tasks. Our TH2T is a two-stage strategy, with tailored dataset construction for each phase, followed by sequential fine-tuning. 1) During the first stage, we use the GSM8K and MATH training datasets to obtain and , respectively. The former is generally considered simple, whereas the latter is regarded as difficult and is explicitly categorized into difficulty levels ranging from 1 to 5, and we pick samples between lv.3 and lv.5. After filtering out errors, and are balanced at a 1:1 ratio. After injecting difficulty hypnosis , we construct of 6.4K samples. 2) In Stage-2, are injected with redundancy hypnosis and looping hypnosis , forming a final training dataset. We observe that the meticulous composition ratio of the constructed dataset does not significantly affect the final performance. Hence, we combine and in a ratio of 1:1. Meanwhile, the redundancy hypnosis and looping hypnosis are also injected into in a ratio of 1:1. The overall composition is set as 2:1:1.
Baselines
In our main experiments, we compare TH2T with two prompt-based methods and two fine-tuning methods.
D-Prompt directly promotes LRM to regulate its output length based on self-judgment of task difficulty based on Difficulty Reminder in Prompt, as discussed in our Preliminary Part.
NoThinking bypasses the thinking process via prompting for all difficulty-level questions.
Tokenskip constructs a dataset where less important tokens are pruned and fine-tuned to enable the model to skip the redundant tokens in inference, regardless of task difficulty.
CoT-Valve enables dynamic control of CoT length by identifying and leveraging a controllable direction in parameter space, reducing inference cost by compressing reasoning paths based on task difficulty.
We replicate the first two baselines and report the relative performance improvement of the latter two from the raw papers due to variability of the model and implementation methods.
Benchmarks
Metrics
Main results are assessed on three metrics: accuracy, average number of output tokens, and inference latency per sample. To substantiate the research objectives, we incorporate the success rate of the difficulty awareness. We also report the redundancy cognition optimization, which includes the reduction of redundant structures (e.g., “Wait”, “Alternatively”, or “Check”), as well as the reflection looping ratio.
Implementation Details
We adopt LoRA, a widespread LLM fine-tuning approach, to train our models with a rank , scaling parameter , and learning rate . All experiments are conducted on 2NVIDIA H800 GPU (80GB). For the decoding strategy, we employ greedy decoding, and the maximum generation length equals 8192 (except 16,384 for AIME2024). The random seed is fixed to 42.
@@
More Experimental Results
Results on Llama Model
@@
Acc and Length on Detailed Difficulty Levels
We report a breakdown of the accuracy and average token length distribution across the complete spectrum of difficulty levels on 4 tasks. The comparison is made between original R1-Ditstilled-Qwen–7B and TH2T-7B. Each question in the MATH and OmniMath datasets is labeled with a difficulty level, ranging from 1 to 5 and from 1 to 8, respectively. The higher the numerical level, the more difficult the question. Meanwhile, all questions in GSM8K and AIME2024 are considered easy and very difficult, respectively. Results are illustrated in Fig.7, and we obtain the following observations: 1) Both models generate progressively longer and more elaborate responses as the difficulty level increases. 2) across all difficulty levels in all tasks, TH2T exhibits significant length reduction, as well as stable performance maintenance. 3) The lower the difficulty level, the more significant the reduction in length.
Dynamics of Accuracy and Token Length
We report the dynamics of accuracy and average token length during the training steps in Fig. 8, and obtain two interesting findings: 1) As training steps increase, accuracy initially declines, then returns to the original level, with occasional positive gains during the process. 2) For hard tasks like MATH, the response length exhibits a reverse increase after exceeding a certain number of steps. This inspires us to set the training steps to 6400 to prevent model degradation.
Visualization of the Effect on Token Length
@@
More Explanation on Motivation
@@
Despite incurring minimal costs, basic LLM can still replicate the majority of correct responses (89.9%) of LRM’s on the easy tasks.
@@
These raise a reasonable speculation: LRMs fall short of perceiving task complexity, since massive inference resources on easy ones are superfluous and lead to limited performance gain. As discussed in Results on Redundancy Cognition of the main paper, we observe severe redundant structures in LRM’s responses to hard tasks, including unnecessary reflection along with a remarkably high proportion of reasoning loops in incorrect answers. These findings lead to the two objectives of this paper:
@@
Difficulty level evaluation
@@
Limitations and Future Work
Our current approach primarily centers on standard fine-tuning using our generated heterogeneous dataset. While this method effectively enhances model performance within our present framework, it may not fully leverage the potential for optimized efficiency. Advanced reinforcement learning (RL)-based training strategies may offer a promising avenue for further performance improvements. However, RL can be very unstable and computationally intensive. Due to computational constraints, we defer the exploration of sophisticated RL techniques to future research.
We also recognize that theoretically, models of different series and parameter scales have distinct capability boundaries, leading to varying levels of task difficulty perception. In this phase, we introduce a common difficulty evaluation standard applicable to models at different scales. We plan to explore difficulty definition methods tailored to models of variant capability ranges in the future, aiming to advance more autonomous difficulty adaptation and perception strategies. One possible solution is to depend on the correctness ratio during sampling. However, a significant limitation of this method lies in its substantially increased computational overhead.
Case Study

### v3

@@ 6 Conclusion
This paper presents TH2T, a novel two-stage fine-tuning strategy for large reasoning models (LRMs) that effectively addresses the challenge of overthinking. Our empirical analysis indicates that current LRMs are largely limited to recognizing task difficulty, exhibiting a necessity for aligning with human System 1 and System 2 thinking patterns. Motivated by this, TH2T progressively enables native difficulty and redundancy cognition based on a hybrid dataset and proposed self-hypnosis mechanism. To the best of our knowledge, we are the first to inject difficulty/redundancy hypnosis as built-in triggers for global, prospective, or local, retrospective interventional signals. Experiments demonstrate that TH2T significantly reduces inference length without compromising performance, exhibiting clear sensitivity to difficulty and redundancy.
Limitations
In real-world cognitive settings, difficulty exists on a continuum, unlike the 2- or 3-level discretization employed in this study. However, to maintain research focus and interpretability, we adopt this simplified cognitive modeling approach. Also, our current approach primarily focuses on standard fine-tuning using the generated hybrid dataset. While this method is effective within our current framework, it may not fully exploit the potential for optimized efficiency. Reinforcement learning presents a promising avenue for further performance gains, although it can be unstable and computationally demanding. Finally, we acknowledge that, in theory, models of parameter scales exhibit distinct capability boundaries, which result in varying levels of task difficulty perception. At this stage, we apply a unified difficulty evaluation standard across models of different scales. Exploring the relationship between a model’s cognition and its capability boundaries is a promising research direction, potentially by dynamic correctness ratio during sampling, despite its significant computational overhead.
Appendices
@@
Appendix A More Experimental Results
A.1 Results on Llama Model
@@
A.2 Ablation on Training Steps.
We examine the dynamics of accuracy and average token length during the training process (in Fig. 7), and obtain two interesting findings: 1) As training steps increase, accuracy initially declines, then returns to the original level, with occasional positive gains during the process. 2) For hard tasks like MATH, the response length exhibits a reverse increase after exceeding a certain number of steps. This inspires us to set the training steps to 6400 to prevent model degradation.
A.3 Visualization of the Effect on Token Length
@@
Appendix B More Experimental Details
B.1 Dataset Construction
The datasets used in this paper incorporate difficulty level labeling as a key consideration, so we mainly use mathematical reasoning tasks. Our TH2T is a two-stage strategy, with tailored dataset construction for each phase, followed by sequential fine-tuning. 1) During the first stage, we use the GSM8K and MATH training datasets to obtain and , respectively. The former is generally considered simple, whereas the latter is regarded as difficult and is explicitly categorized into difficulty levels ranging from 1 to 5, and we pick samples between lv.3 and lv.5. After filtering out errors, and are balanced at a 1:1 ratio. After injecting difficulty hypnosis , we construct of 6.4K samples. 2) In Stage-2, are injected with redundancy hypnosis and looping hypnosis , forming a final training dataset. We observe that the meticulous composition ratio of the constructed dataset does not significantly affect the final performance. Hence, we combine and in a ratio of 1:1. Meanwhile, the redundancy hypnosis and looping hypnosis are also injected into in a ratio of 1:1. The overall composition is set as 2:1:1, yielding samples in total.
B.2 Benchmarks
B.3 Metrics
Main results are assessed on accuracy and the average number of output tokens. To substantiate the research objectives, we incorporate the success rate of the difficulty awareness. We also report the redundancy cognition optimization, which includes the reduction of redundant structures (e.g., “Wait”, “Alternatively”, or “Check”), as well as the reflection looping ratio. To examine whether our method instills a genuine sense of difficulty cognition, we investigate the model’s internal state, i.e., the confidence value of the first generated token.
B.4 Implementation Details
We adopt LoRA, a widespread LLM fine-tuning approach, to train our models with a rank , scaling parameter , and learning rate . All experiments are conducted on 4NVIDIA H800 GPU (80GB). For the decoding strategy, we employ greedy decoding. The maximum generation length is set to 8192 for GSM8K and MATH500, and 16,384 for AIME2024, OmniMath, and GPQA. The random seed is fixed to 42.
Appendix C More Explanation on Motivation
@@
Despite incurring minimal costs, basic LLM can still replicate the majority of correct responses (89.9%) of LRM’s on the easy tasks, which is why we claimed that can replicate the majority correct responses of on in Sec.4.1.
@@
These raise a reasonable speculation: LRMs fall short of perceiving task complexity, since massive inference resources on easy ones are superfluous and lead to limited performance gain. As discussed in Results on Redundancy Cognition of the main paper, we observe severe redundant structures in LRM’s responses to hard tasks, including superfluous reflection along with a remarkably high proportion of reasoning loops in incorrect answers. These findings lead to the two objectives of this paper:
@@ Appendix D Prompts
D.1 Difficulty level evaluation
@@ Appendix E Design of Determinator
Appendix F Case Study